OpenStudy (anonymous):

If point C is the midpoint of AB , what is the coordinate of point B? C=(1,4), A=(5,5)

OpenStudy (tasreeb):

c=x1+x2/2 1=5+x/2 2=5+x x=-3, 4=5+x/2 8=5+x x=3 B(3,-3)